The Leos invested part of $12,000 at 4% simple interest per year and the rest at 6% simple interest per year. The total interest for one year was at most $640. Find the least amount that could have been invested at 4%.
Answer by checkley77(12844)

.06x+.04(12,000-x)<=640
.06x+480-.04x<=640
.02x<=640-480
.02x<=160
x<=160/.02
x=$8,000 amount invested @ 6%.
12,000-8,000=$4,000 amount invested @ 4%
Proof:
.06*8,000+.04*4,000<=640
480+160<=640
640<=640